# A Bregman ADMM for Bethe variational problem

### v2

@@ 1.1 Graphical model and belief propagation
Graphical model [34] is a powerful framework for representing multivariate distributions, with a graph encoding the dependency structure among variables. It has applications in many fields including scientific computing [59], machine learning [6] and statistical physics [58]. A key challenge in graphical models is probabilistic inference, which we describe in the next paragraph.
@@
where $x_{k}$ is the state of node $k$ with $r$ possible values, $\Psi_{ij}(x_{i},x_{j})$ and $\Psi_{k}(x_{k})$ are potential functions, and $Z$ is a normalization constant. The goal of probabilistic inference is to determine the marginal distribution of a single random variable $x_{k}$. While the marginal distribution can be computed directly by summing over all other variables, this approach is generally NP-hard [16] due to the exponential growth in the state space as $n$ increases.
Belief propagation. In 1988, Pearl introduced belief propagation (BP) [57] as an approximate method to solve the probabilistic inference problem. For each edge $ij$ (In this paper, we assume $i<j$ when referring to $ij\in\mathcal{E}$.) in $\mathcal{E}$, BP defines message variables $m_{i\rightarrow j}(x_{j})$ and $m_{j\rightarrow i}(x_{i})$, which are updated as follows:
@@
where $\hat{z}_{k}$ and $\bar{z}_{i,j}$ are normalization constants ensuring that $b_{k}$ and $b_{i,j}$ are valid probability distributions. The approximations $b_{k}$ and $b_{i,j}$, often referred to as beliefs, provide estimates of the marginal distributions of the joint probability $P$. For tree-structured graphs, BP is guaranteed to converge to the exact marginal distributions in a finite number of iterations. However, for graphs with cycles, convergence is not guaranteed [54]. Nevertheless, BP often performs remarkably well as a heuristic in practice and has been successfully applied in various domains [1, 47, 53].
@@ 1.2 Bethe variational problem
The practical success of BP has inspired extensive theoretical investigations. Yedidia, Freeman, and Weiss achieved a key breakthrough in their seminal works [83, 84, 82], where they demonstrated that the fixed points of BP correspond to the critical points of the following Bethe variational problem (BVP)[73]:
[…]
@@
where $b_{k}$ and $b_{ij}$ are the beliefs satisfying normalization and marginalization constraints. The objective function in (6), known as the Bethe free energy (BFE), was introduced by Hans Bethe in 1935 [8]. It serves as a powerful approximation tool in statistical physics [46], simplifying the computation of free energy in systems where exact calculations are computationally infeasible. Alternatively, the BFE can be interpreted as the Kullback-Leibler (KL) divergence [32] between a tree-like probability distribution
@@
However, one important limitation of the earlier works by Yedidia et al. [83, 84, 82] is that they considered only the variables $b_{ij}$ and $b_{k}$ with strictly positive entries, thereby ignoring boundary points in the feasible set of (6). If the minimizer of (6) contains zero entries, it cannot serve as a fixed point for BP. However, boundary points cannot be disregarded when solving (6), as the non-convexity makes it challenging to ensure that local minima will not have zero entries. Furthermore, zero entries pose significant challenges for optimization algorithms because the objective function is not Lipschitz-continuous on the boundary. This observation motivates the following question:
@@ 1.3 Algorithms for Bethe variational problem
Gradient descent method. Welling and Teh [78] proposed a gradient descent (GD) method to find a stationary point of (BVP) for $binary$ MRFs. The key insight is that for $r=2$, problem (BVP) can be reformulated as an unconstrained optimization problem with $n$ variables, allowing GD to be applied. Shin refined this approach in [66, 67] to get a polynomial time complexity bound for graphs with maximum degree $\mathcal{O}(\log n)$. Additionally, Leisenberger et al. [36] extended this idea using a projected quasi-Newton method to exploit second-order information. For binary pairwise MRFs with $submodular$ cost functions, global minimizers of (BVP) have been explored using GD [77]. Despite its theoretical soundness, the GD method is limited to binary networks. For $r>2$, it is unclear how to reformulate the problem as unconstrained optimization, rendering GD inapplicable in such cases.
Double-loop algorithm. Yuille [86, 87] introduced a double-loop algorithm called the concave-convex procedure (CCCP) to solve (BVP) for any $r\geq 2$. The key observation is that the objective function of (BVP) is a difference of convex (DC) functions, enabling the application of DC programming techniques [26]. In each outer iteration of CCCP, the concave part is linearized to create a majorization. The obtained convex problem is then solved using a block coordinate minimization (BCM) method applied to its dual problem. Although CCCP can handle any $r\geq 2$, its efficiency is hindered by the iterative nature of BCM, as solving subproblems to high accuracy can be computationally expensive. Furthermore, while CCCP guarantees monotonic decreases in the objective function, the convergence of the variables remains unresolved, motivating the need for algorithms with more rigorous theoretical guarantees. The strengths and weaknesses of classical algorithms for (BVP) are summarized in Table 1.
@@ 1.4 Quantum Bethe variational problem
Akin to classical graphical models, Leifer and Poulin introduced quantum graphical models in [35], with applications in quantum error correction [62], simulations of many-body quantum systems [52], and approximations of ground-state energy [63]. Hastings [24] further proposed quantum belief propagation (QBP) as a quantum counterpart to classical belief propagation for solving quantum graphical models. Recently, Zhao et al. [88] explored the connection between QBP and the quantum Bethe variational problem (5.1).
A related class of optimization problems involves semidefinite programming (SDP) relaxations of quantum many-body problems via quantum embedding theories [69, 21, 31, 29, 30]. These models are similar to (5.1) in that they approximate a many-body quantum system by clustering accurate local densities and combining them using local consistency constraints. In particular, Lin and Lindsey [42] proposed tightening convex relaxations by intersecting local consistency constraints with the global semidefinite constraint. More recently, Khoo and Lindsey [28] developed an ADMM-type method to efficiently solve these SDP problems by exploiting the parallelizability of the subproblems. However, the method proposed there does not readily work with entropy, and also there is no convergence guarantee. Compared to the model in [42], (5.1) does not impose a global semidefinite constraint but instead incorporates an additional entropy function. In this work, we focus on (5.1), though we believe extending our approach to solving SDP problems in quantum embedding theories is a promising direction for future research.
A separate class of convex optimization problems stems from two-electron reduced density matrix (2-RDM) theory [49, 50, 13, 51, 89, 38, 2, 56, 17]. This approach approximates the many-body density function using low-order moments [33]. The formulation of 2-RDM optimization problems differs significantly from both (5.1) and quantum embedding theories, so we do not pursue it in this work.
@@ 1.7 Notations and definitions
[…]
    $\mathcal{Y}$ ::=:= $\left\{{\bf Q}\in\mathcal{H}_{2}:\ \forall ij\in\mathcal{E},\ Q_{ij}\in\mathbb{R}^{r\times r}_{+},\ \left\langle{\bf 1}_{r\times r},Q_{ij}\right\rangle=1\right\}.$ (12)
@@
    $\mathcal{F}({\bf q}):=\left\langle{\bf c},{\bf q}\right\rangle-\sum_{k\in[n]}(d_{k}-1)\left\langle q_{k},\log q_{k}\right\rangle,\mathcal{G}({\bf Q}):=\left\langle{\bf C},{\bf Q}\right\rangle+\sum_{ij\in\mathcal{E}}\left\langle\log Q_{ij},Q_{ij}\right\rangle.$ (14)
@@
  • For any strictly convex function $\phi,$ the Bregman divergence $B_{\phi}(\cdot,\cdot)$ [12] is defined as follows:
@@
    ${\rm Im}(\mathcal{A}\circ\mathcal{P}_{\mathcal{X}})\subset{\rm Im}(\mathcal{B}\circ\mathcal{P}_{\mathcal{Y}})={\rm Im}(\hat{\mathcal{P}}).$ (23)
@@
    $\xi_{r,\tau}$ ::=:= $\frac{\sqrt{|\mathcal{E}|}r\left(-2\log\tau+2\log 2r+5M\right)}{\sqrt{\gamma_{r}}},$ (24)
    $\hat{\xi}_{r,\tau}$ ::=:= $2nM+2(2|\mathcal{E}|-n)\log r+2\xi_{r,\tau}\sqrt{\xi_{r,\tau}^{2}+4|\mathcal{E}|(M+\log r)}$ (25)
@@
    $\tilde{\xi}_{r,\tau}$ ::=:= $4\hat{\xi}_{r,\tau}+2\sqrt{\xi_{r,\tau}^{2}+4|\mathcal{E}|(M+\log r)}.$ (26)
@@
[…]
  • Consider linear space ${\bf X},$ a function $f:{\bf X}\rightarrow\mathbb{R}\cup\{+\infty\}$ is said to have the Kurdyka-Łojasiewicz (KL) property [4] at $x^{*}\in{\rm dom}(\partial f)$ if there exists $\eta\in(0,+\infty],$ a neighborhood $U$ of $x^{*}$ and a continuous concave function $\varphi:[0,\eta)\rightarrow\mathbb{R}_{+}$ such that:
@@
    Functions satisfying the KL property including semi-algebraic, subanalytic and log-exp functions [3, 5].
@@ 2 Local minima of (BVP)
Yedidia et al. provided a negative answer to Question 1.1 in [85, Theorem 9]. However, their proof only addressed the case $r=2$ and claimed that the argument could be easily extended to the general case. While their contribution is significant, we argue that the case $r>2$ is much more intricate than $r=2$. For $r=2$, problem (BVP) can be reformulated as an unconstrained optimization problem with $n$ variables [78], whose gradient has an explicit form, which significantly simplifies the perturbation analysis. In contrast, for $r\geq 3$, the constraints cannot be eliminated, making the corresponding perturbation analysis considerably more complex.
@@
The main advantage of solving (BVPτ) lies in the constraint ${\bf q}\geq\tau$, which prevents algorithmic iterations from producing zero entries. This constraint plays a crucial role in the convergence analysis of our algorithms. Since finding local minima for non-convex optimization problems is generally NP-hard [55], we establish the following equivalence result regarding the stationarity of (BVP) and (BVPτ):
@@ 2.2 Properties of entropy regularized optimal transport
Before we prove Theorem 2.1 and Proposition 2.2, we first prove several lemmas about entropy regularized optimal transport (OT) problem [15].
@@
From Lemma 2 in [15],we know that the optimal solution $X^{*}$ of (31) can be written in the form (32) for some $\lambda,\mu\in\mathbb{R}^{r}.$ Note that the Lagrangian multipliers $\lambda,\mu$ are not unique because for any $\alpha\in\mathbb{R},$ $\lambda^{\prime}:=\lambda-\alpha{\bf 1}_{r},$ $\mu^{\prime}:=\mu+\alpha{\bf 1}_{r}$ also satisfy the equation (32). Thus, without loss of generality, we assume $\max_{i\in[r]}\{\lambda(i)\}=\max_{i\in[r]}\{\mu(i)\}.$ Let $a,b\in[r]$ such that $\lambda(a):=\max_{i\in[r]}\{\lambda(i)\},$ $\mu(b):=\max_{i\in[r]}\{\mu(i)\}.$ From (32), we have that
@@ 2.3 Equivalent formulation of (BVP)
From Proposition 4.6 of [60], we know that $L_{ij}:\triangle_{r}\times\triangle_{r}\rightarrow\mathbb{R}$ is a convex function and for any $(q_{i},q_{j})\in\operatorname{ri}(\triangle_{r})\times\operatorname{ri}(\triangle_{r}),$ it is differentiable with gradient given as follows:
@@
:=$\mathcal{I}:=\left\{k\in[n]:\ \exists\ i\in[r]\ {\rm s.t.}\ q^{*}_{k}(i)<\sigma\right\}.$
@@
Because $\sigma\in(0,1/r),$ for any $k\in\mathcal{I},$ $11\leq|J_{k}|\leq r-$ 1 . Define the direction ${\bf v}=(v_{k})_{k\in[n]}\in\mathcal{H}_{1}$ such that
@@
[…]
@@
Because $({\bf q}^{*},{\bf Q}^{*})$ is a local minimizer of (BVP), we have that ${\bf Q}^{*}=(Q^{*}_{ij})_{ij\in\mathcal{E}}$ are global minimizers of the convex problems (40) with given marginals ${\bf q}^{*}.$ Assume on the contrary that ${\bf q}^{*}\geq\sigma$ doesn’t hold. From Lemma 2.4, there exists a descent feasible direction ${\bf v}$ at ${\bf q}^{*}$ for (42). Therefore, ${\bf q}^{*}$ is not a local minimizer of (42). From the continuity of the optimal solution of the entropy OT problem (45) (see [19, Theorem 3.11]), we know that $({\bf q}^{*},{\bf Q}^{*})$ is also not a local minimizer of (BVP). We get a contradiction. Therefore, ${\bf q}^{*}\geq\sigma.$ From (35) in Lemma 2.3, we have that ${\bf Q}^{*}\geq\sigma^{2}/(r^{2}\exp[4M]).$

### v3

@@ 1.1 Graphical model and belief propagation
Graphical model [lauritzen1996graphical] is a powerful framework for representing multivariate distributions, with a graph encoding the dependency structure among variables. It has applications in many fields including scientific computing [perry2018message], machine learning [bach2004learning] and statistical physics [pelizzola2005cluster]. A key challenge in graphical models is probabilistic inference, which we describe in the next paragraph.
@@
where $x_{k}$ is the state of node $k$ with $r$ possible values, $\Psi_{ij}(x_{i},x_{j})$ and $\Psi_{k}(x_{k})$ are potential functions, and $Z$ is a normalization constant. The goal of probabilistic inference is to determine the marginal distribution of a single random variable $x_{k}$. While the marginal distribution can be computed directly by summing over all other variables, this approach is generally NP-hard [dagum1993approximating] due to the exponential growth in the state space as $n$ increases.
Belief propagation. In 1988, Pearl introduced belief propagation (BP) [pearl2014probabilistic] as an approximate method to solve the probabilistic inference problem. For each edge $ij$ (In this paper, we assume $i<j$ when referring to $ij\in\mathcal{E}$.) in $\mathcal{E}$, BP defines message variables $m_{i\rightarrow j}(x_{j})$ and $m_{j\rightarrow i}(x_{i})$, which are updated as follows:
@@
where $\hat{z}_{k}$ and $\bar{z}_{i,j}$ are normalization constants ensuring that $b_{k}$ and $b_{i,j}$ are valid probability distributions. The approximations $b_{k}$ and $b_{i,j}$, often referred to as beliefs, provide estimates of the marginal distributions of the joint probability $P$. For tree-structured graphs, BP is guaranteed to converge to the exact marginal distributions in a finite number of iterations. However, for graphs with cycles, convergence is not guaranteed [murphy2013loopy]. Nevertheless, BP often performs remarkably well as a heuristic in practice and has been successfully applied in various domains [abbe2018proof, makur2024robustness, mridula2011combining].
@@ 1.2 Bethe variational problem
The practical success of BP has inspired extensive theoretical investigations. Yedidia, Freeman, and Weiss achieved a key breakthrough in their seminal works [yedidia2001bethe, yedidia2003understanding, yedidia2000generalized], where they demonstrated that the fixed points of BP correspond to the critical points of the following Bethe variational problem (BVP)[wainwright2008graphical]:
$\min\Bigg\{$ $\sum_{ij\in\mathcal{E}}\sum_{x_{i},x_{j}\in[r]}b_{ij}(x_{i},x_{j})\ln\frac{b_{ij}(x_{i},x_{j})}{\Psi_{ij}(x_{i},x_{j})b_{i}(x_{i})b_{j}(x_{j})}+\sum_{k\in V}\sum_{x_{k}\in[r]}b_{k}(x_{k})\ln\frac{b_{k}(x_{k})}{\Psi_{k}(x_{k})}:$ (6)
@@
where $b_{k}$ and $b_{ij}$ are the beliefs satisfying normalization and marginalization constraints. The objective function in (6), known as the Bethe free energy (BFE), was introduced by Hans Bethe in 1935 [bethe1935statistical]. It serves as a powerful approximation tool in statistical physics [mackay2003information], simplifying the computation of free energy in systems where exact calculations are computationally infeasible. Alternatively, the BFE can be interpreted as the Kullback-Leibler (KL) divergence [kullback1951information] between a tree-like probability distribution
@@
However, one important limitation of the earlier works by Yedidia et al. [yedidia2001bethe, yedidia2003understanding, yedidia2000generalized] is that they considered only the variables $b_{ij}$ and $b_{k}$ with strictly positive entries, thereby ignoring boundary points in the feasible set of (6). If the minimizer of (6) contains zero entries, it cannot serve as a fixed point for BP. However, boundary points cannot be disregarded when solving (6), as the non-convexity makes it challenging to ensure that local minima will not have zero entries. Furthermore, zero entries pose significant challenges for optimization algorithms because the objective function is not Lipschitz-continuous on the boundary. This observation motivates the following question:
@@ 1.3 Algorithms for Bethe variational problem
Gradient descent method. Welling and Teh [welling2013belief] proposed a gradient descent (GD) method to find a stationary point of (BVP) for binarybinary MRFs. The key insight is that for $r=2$, problem (BVP) can be reformulated as an unconstrained optimization problem with $n$ variables, allowing GD to be applied. Shin refined this approach in [shin2012complexity, shin2014complexity] to get a polynomial time complexity bound for graphs with maximum degree $\mathcal{O}(\log n)$. Additionally, Leisenberger et al. [leisenberger2024convexity] extended this idea using a projected quasi-Newton method to exploit second-order information. For binary pairwise MRFs with submodularsubmodular cost functions, global minimizers of (BVP) have been explored using GD [weller2013bethe]. Despite its theoretical soundness, the GD method is limited to binary networks. For $r>2$, it is unclear how to reformulate the problem as unconstrained optimization, rendering GD inapplicable in such cases.
Double-loop algorithm. Yuille [yuille2001double, yuille2002cccp] introduced a double-loop algorithm called the concave-convex procedure (CCCP) to solve (BVP) for any $r\geq 2$. The key observation is that the objective function of (BVP) is a difference of convex (DC) functions, enabling the application of DC programming techniques [horst1999dc]. In each outer iteration of CCCP, the concave part is linearized to create a majorization. The obtained convex problem is then solved using a block coordinate minimization (BCM) method applied to its dual problem. Although CCCP can handle any $r\geq 2$, its efficiency is hindered by the iterative nature of BCM, as solving subproblems to high accuracy can be computationally expensive. Furthermore, while CCCP guarantees monotonic decreases in the objective function, the convergence of the variables remains unresolved, motivating the need for algorithms with more rigorous theoretical guarantees. The strengths and weaknesses of classical algorithms for (BVP) are summarized in Table 1.
@@ 1.4 Quantum Bethe variational problem
Akin to classical graphical models, Leifer and Poulin introduced quantum graphical models in [leifer2008quantum], with applications in quantum error correction [poulin2008belief], simulations of many-body quantum systems [mezard2007constraint], and approximations of ground-state energy [poulin2011markov]. Hastings [hastings2007quantum] further proposed quantum belief propagation (QBP) as a quantum counterpart to classical belief propagation for solving quantum graphical models. Recently, Zhao et al. [zhao2024quantum] explored the connection between QBP and the quantum Bethe variational problem (5.1).
A related class of optimization problems involves semidefinite programming (SDP) relaxations of quantum many-body problems via quantum embedding theories [sun2016quantum, georges1996dynamical, kotliar2006electronic, knizia2012density, knizia2013density]. These models are similar to (5.1) in that they approximate a many-body quantum system by clustering accurate local densities and combining them using local consistency constraints. In particular, Lin and Lindsey [lin2022variational] proposed tightening convex relaxations by intersecting local consistency constraints with the global semidefinite constraint. More recently, Khoo and Lindsey [khoo2024scalable] developed an ADMM-type method to efficiently solve these SDP problems by exploiting the parallelizability of the subproblems. However, the method proposed there does not readily work with entropy, and also there is no convergence guarantee. Compared to the model in [lin2022variational], (5.1) does not impose a global semidefinite constraint but instead incorporates an additional entropy function. In this work, we focus on (5.1), though we believe extending our approach to solving SDP problems in quantum embedding theories is a promising direction for future research.
A separate class of convex optimization problems stems from two-electron reduced density matrix (2-RDM) theory [mazziotti1998contracted, mazziotti2004realization, cances2006electronic, mazziotti2012structure, zhao2004reduced, li2018semismooth, anderson2013second, nakata2001variational, deprince2010exploiting]. This approach approximates the many-body density function using low-order moments [lasserre2001global]. The formulation of 2-RDM optimization problems differs significantly from both (5.1) and quantum embedding theories, so we do not pursue it in this work.
@@ 1.7 Notations and definitions
    $\mathcal{X}$ $:=$ $\left\{{\bf q}\in\mathcal{H}_{1}:\ \forall k\in[n],\ q_{k}\in\mathbb{R}^{r}_{+},\ {\bf 1}_{r}^{\top}q_{k}=1\right\},$ (11)
    $\mathcal{Y}$ $:=$ $\left\{{\bf Q}\in\mathcal{H}_{2}:\ \forall ij\in\mathcal{E},\ Q_{ij}\in\mathbb{R}^{r\times r}_{+},\ \left\langle{\bf 1}_{r\times r},Q_{ij}\right\rangle=1\right\}.$ (12)
@@
    $\hskip-19.91692pt\mathcal{F}({\bf q}):=\left\langle{\bf c},{\bf q}\right\rangle-\sum_{k\in[n]}(d_{k}-1)\left\langle q_{k},\log q_{k}\right\rangle,\mathcal{G}({\bf Q}):=\left\langle{\bf C},{\bf Q}\right\rangle+\sum_{ij\in\mathcal{E}}\left\langle\log Q_{ij},Q_{ij}\right\rangle.$ (14)
@@
  • For any strictly convex function $\phi,$ the Bregman divergence $B_{\phi}(\cdot,\cdot)$ [bregman1967relaxation] is defined as follows:
@@
[…]
@@
    $\xi_{r,\tau}$ $:=$ $\frac{\sqrt{|\mathcal{E}|}r\left(-2\log\tau+2\log 2r+5M\right)}{\sqrt{\gamma_{r}}},$ (24)
    $\hat{\xi}_{r,\tau}$ $:=$ $2nM+2(2|\mathcal{E}|-n)\log r+2\xi_{r,\tau}\sqrt{\xi_{r,\tau}^{2}+4|\mathcal{E}|(M+\log r)}$ (25)
@@
    $\tilde{\xi}_{r,\tau}$ $:=$ $4\hat{\xi}_{r,\tau}+2\sqrt{\xi_{r,\tau}^{2}+4|\mathcal{E}|(M+\log r)}.$ (26)
@@
    $\triangle_{r}:=\left\{x\in\mathbb{R}^{r}:\ {\bf 1}_{r}^{\top}x=1,\ x\geq 0\right\}.$ (27)
  • Consider linear space $\bf X$}, a function $f:{\bf X}\rightarrow\mathbb{R}\cup\{+\infty\}$ is said to have the Kurdyka-Łojasiewicz (KL) property [attouch2010proximal] at $x^{*}\in{\rm dom}(\partial f)$ if there exists $\eta\in(0,+\infty],$ a neighborhood $U$ of $x^{*}$ and a continuous concave function $\varphi:[0,\eta)\rightarrow\mathbb{R}_{+}$ such that:
@@
    Functions satisfying the KL property including semi-algebraic, subanalytic and log-exp functions [attouch2009convergence, attouch2013convergence].
@@ 2 Local minima of (BVP)
Yedidia et al. provided a negative answer to Question 1.1 in [yedidia2005constructing, Theorem 9]. However, their proof only addressed the case $r=2$ and claimed that the argument could be easily extended to the general case. While their contribution is significant, we argue that the case $r>2$ is much more intricate than $r=2$. For $r=2$, problem (BVP) can be reformulated as an unconstrained optimization problem with $n$ variables [welling2013belief], whose gradient has an explicit form, which significantly simplifies the perturbation analysis. In contrast, for $r\geq 3$, the constraints cannot be eliminated, making the corresponding perturbation analysis considerably more complex.
@@
The main advantage of solving (BVPτ) lies in the constraint ${\bf q}\geq\tau$, which prevents algorithmic iterations from producing zero entries. This constraint plays a crucial role in the convergence analysis of our algorithms. Since finding local minima for non-convex optimization problems is generally NP-hard [quadNPhard], we establish the following equivalence result regarding the stationarity of (BVP) and (BVPτ):
@@ 2.2 Properties of entropy regularized optimal transport
Before we prove Theorem 2.1 and Proposition 2.2, we first prove several lemmas about entropy regularized optimal transport (OT) problem [cuturi2013sinkhorn].
@@
From Lemma 2 in [cuturi2013sinkhorn],we know that the optimal solution $X^{*}$ of (31) can be written in the form (32) for some $\lambda,\mu\in\mathbb{R}^{r}.$ Note that the Lagrangian multipliers $\lambda,\mu$ are not unique because for any $\alpha\in\mathbb{R},$ $\lambda^{\prime}:=\lambda-\alpha{\bf 1}_{r},$ $\mu^{\prime}:=\mu+\alpha{\bf 1}_{r}$ also satisfy the equation (32). Thus, without loss of generality, we assume $\max_{i\in[r]}\{\lambda(i)\}=\max_{i\in[r]}\{\mu(i)\}.$ Let $a,b\in[r]$ such that $\lambda(a):=\max_{i\in[r]}\{\lambda(i)\},$ $\mu(b):=\max_{i\in[r]}\{\mu(i)\}.$ From (32), we have that
@@ 2.3 Equivalent formulation of (BVP)
From Proposition 4.6 of [peyre2019computational], we know that $L_{ij}:\triangle_{r}\times\triangle_{r}\rightarrow\mathbb{R}$ is a convex function and for any $(q_{i},q_{j})\in\operatorname{ri}(\triangle_{r})\times\operatorname{ri}(\triangle_{r}),$ it is differentiable with gradient given as follows:
@@
$\mathcal{I}:=\left\{k\in[n]:\ \exists\ i\in[r]\ {\rm s.t.}\ q^{*}_{k}(i)<\sigma\right\}.$
@@
Because $\sigma\in(0,1/r),$ for any $k\in\mathcal{I},$ $1\leq|J_{k}|\leq r-1.$ Define the direction ${\bf v}=(v_{k})_{k\in[n]}\in\mathcal{H}_{1}$ such that
@@
Therefore, $\bf v$} is a feasible direction of (42) that satisfies (i). Next, we proceed to prove (ii). For any $t\in(0,\delta),\ \penalty 10000\thinspace i\kern-0.20004pt\penalty 10000\thinspace j\in\mathcal{E},$ consider the following entropy OT problems:
@@
Because $({\bf q}^{*},{\bf Q}^{*})$ is a local minimizer of (BVP), we have that ${\bf Q}^{*}=(Q^{*}_{ij})_{ij\in\mathcal{E}}$ are global minimizers of the convex problems (40) with given marginals $\bf q$$^{*}$. Assume on the contrary that ${\bf q}^{*}\geq\sigma$ doesn’t hold. From Lemma 2.4, there exists a descent feasible direction $\bf v$} at $\bf q$$^{*}$ for (42). Therefore, $\bf q$$^{*}$ is not a local minimizer of (42). From the continuity of the optimal solution of the entropy OT problem (45) (see [eckstein2022quantitative, Theorem 3.11]), we know that $({\bf q}^{*},{\bf Q}^{*})$ is also not a local minimizer of (BVP). We get a contradiction. Therefore, ${\bf q}^{*}\geq\sigma.$ From (35) in Lemma 2.3, we have that ${\bf Q}^{*}\geq\sigma^{2}/(r^{2}\exp[4M]).$